OpenStudy (blaque678):

(5,-1),(9,3),(0,-6)(-6,-12)(9 1/2,?) what's the next number?

OpenStudy (anonymous):

let the points (a,b) and (c,d). the number d is calculated by this b+c-a [here only] so the number is, -12+[9(1/2)-(-6)]=-12+[19/2+6]=-12+31/2=7/2=3(1/2)=3 1/2